# Gradient Networks

### v2

@@ 1 Introduction
Deep neural networks are prized for their ability to parameterize and easily learn complicated, high-dimensional functions. Researchers have devoted substantial effort to developing deep neural networks that achieve state of the art performance on numerous tasks spanning computer vision [1], natural language processing [2], and reinforcement learning [3]. These neural networks are commonly unconstrained and effectively parameterize the space of all functions. Many applications instead require learned functions to exhibit specific properties, which necessitates the design of neural networks corresponding to specific function classes – a problem that has seldom been studied. Constraining neural networks to belong to a particular function class not only enhances interpretability, but also leads to theoretical performance guarantees essential for deploying trained models in safety-critical applications.
In particular, neural networks that correspond to gradients of functions hold significant importance across several science and engineering disciplines. For example, physicists often wish to use a finite set of measurements to characterize the gradient field of a potential function, such as the temperature gradient over a surface. Score-based generative models are another application of learning gradient functions, where neural networks are trained to learn $\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x})$, the score function of an unknown probability distribution [4, 5]. These methods have been applied to image [6, 7] and 3D shape [8] generation.
@@
Contributions: We propose gradient networks (GradNets), neural networks for directly parameterizing gradients of arbitrary functions. We adapt GradNets to monotone gradient networks (mGradNets) that correspond to gradients of convex functions. Our proposed neural network architectures directly model the gradient function $\nabla F$ without first learning the underlying potential function $F$. We also analyze the expressivity of our proposed GradNet and mGradNet networks and prove that they are universal approximators of gradients of general and convex functions respectively. We further describe methods for modifying these networks to represent popular subsets of these function classes, including gradients of sums of (convex) ridge functions and their (convexity-preserving) transformations. Our theoretical results translate into improved experimental results over existing works. In our gradient field approximation experiments, we find that our architectures achieve up to 15 dB lower mean squared error (MSE) than existing methods. Additionally, we address the problem of predicting the dynamics of the two-body problem and demonstrate an 11 dB improvement over existing methods.
Paper Organization: We first review existing methods for representing gradients of functions in Sec. 2. We then present gradient networks (GradNets) for learning gradients of arbitrary functions in Sec. 3. In Sec. 4, we describe monotone gradient networks (mGradNets): modified GradNets for representing gradients of convex functions. We analyze the expressivity of our proposed GradNets and mGradNets in Sec. 5 and demonstrate methods for composing GradNets and mGradNets to learn (sub)gradients of of the function classes described in Fig. 1. We present architectural modifications in Sec. 6 that empirically yield efficient parameterizations. Finally, in Sec. 7, we evaluate our proposed models on gradient field and Hamiltonian dynamics learning tasks.
@@ 2 Related Work
Several existing works use standard neural networks to directly model the gradient of a potential function [23, 7, 24, 17]. While these methods demonstrate satisfactory empirical performance, they lack theoretical justification and are often considered heuristic approaches. It is not guaranteed, and in fact highly unlikely, that an arbitrary neural network with matching input and output dimension corresponds to the gradient of a scalar-valued function. More formally, let $\mathcal{F}$ be a set such that each function $F\in\mathcal{F}$ is the gradient of a scalar-valued function. Given an $F\in\mathcal{F}$, standard neural network architectures, e.g., multilayer perceptrons (MLPs) and convolutional neural networks (CNNs), with nonpolynomial activations can be used to approximate $F$. However, these architectures offer no guarantee that the learned function $\widehat{F}$ is itself in $\mathcal{F}$. In contrast, the methods and network architectures discussed in this paper ensure that the approximator $\widehat{F}$ is always a member of $\mathcal{F}$, thereby enhancing interpretability and enabling robust theoretical guarantees in practice.
@@
Rather than directly approximating the gradient of a function using a standard neural network, an alternative approach may be to first parameterize and learn the underlying scalar potential function using a standard neural network. Subsequently differentiating the network with respect to its input yields the desired gradient. However, closely approximating a target function $F$ with an arbitrary approximator $G$ does not guarantee that $\nabla G$ closely approximates $\nabla F$ [25]. Empirical evidence in [26] confirms this fact when the approximating functions are neural networks. Other works directly train the gradient map of a neural network that parameterizes a potential function. For instance, [21] considers the potential $F(\boldsymbol{x})=\frac{1}{2}\|\boldsymbol{x}-f_{\boldsymbol{\theta}}(\boldsymbol{x})\|_{2}^{2}$ with neural network $f_{\boldsymbol{\theta}}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and uses automatic differentiation to obtain a gradient network architecture $\nabla F(\boldsymbol{x})$. Nevertheless, these approaches often exhibit ill-behaved product structures in practice [27] and, by the Runge phenomena [28], optimizing their parameters becomes cumbersome as the input dimension grows [29].
@@
Feedforward neural networks have also been proposed to directly parameterize gradients of specific classes of potential functions, namely potentials $F(\boldsymbol{x})$ expressible as the sum of ridge functions: $F(\boldsymbol{x})=\sum_{i=1}^{n}\psi_{i}(\boldsymbol{w}_{i}^{\top}\boldsymbol{x})$. Such potentials were considered in the fields of experts framework proposed in [30]. The fields of experts framework has inspired several methods for learning image priors that effectively generalize the “transform domain thresholding” approach discussed in [17, 31]. These include works like [32, 33] that correspond to nonconvex potentials and respectively achieve commendable performance for image restoration and magnetic resonance imaging reconstruction. More recently, [34, 35] propose neural networks with elementwise, learnable spline activations to learn gradients of potentials expressible as sums of ridge functions (class G in Fig. 1). The networks can be modified to correspond to gradients of sums of convex ridge functions (class H in Fig. 1) by restricting the splines to be nondecreasing. In this paper, we consider parameterizing the activations as learnable neural networks and discuss how these networks can be adapted to represent (sub)gradients of convex potentials. Furthermore, the class of functions expressible as a sum of convex ridges does not include several common functions including
@@ 3 Gradient Networks (GradNet)
Prop. 0.1 states that (3) is a GradNet if the activation function $\sigma$ has an antiderivative. It hence permits the use of group activations such as softmax. Furthermore, to guarantee existence of an antiderivative, it is sufficient, but not necessary, that $\sigma$ is differentiable and its Jacobian is symmetric everywhere. For example, the elementwise activation $\mathrm{ReLU}(\boldsymbol{x})=\max(0,\boldsymbol{x})$ is continuous and nondifferentiable, but it is the gradient of $\sum_{i}\max(0,\frac{1}{2}\boldsymbol{x}_{i}^{2})$. Architectures of the form in (3), with a continuous elementwise activation function $\sigma(\boldsymbol{x})=\begin{bmatrix}\sigma_{1}(\boldsymbol{x}_{1})\ \dots\ \sigma_{m}(\boldsymbol{x}_{m})\end{bmatrix}^{\top}$, include “transform domain thresholding” methods [17, 31] where the activation is a continuous thresholding operator. Similar architectures appear in [35], which uses linear spline activations, and in [33], which parameterizes each $\sigma_{i}$ as a linear combination of Gaussian radial basis functions. An alternative approach is to use neural networks to parameterize the elementwise activation functions.
@@ 4 Monotone Gradient Networks (mGradNet)
Rem. 3 follows from Def. 1 and Remark 2.2 in [35], which implies monotonicity of $g$. It demonstrates how to construct an mGradNet using a GradNet with known Lipschitz constant. Rem. 3 relates to several works that design learnable, monotone functions of the form $\boldsymbol{x}-f(\boldsymbol{x})$ by restricting $f(\cdot)$ to be nonexpansive [45, 46]. These methods ensure $\boldsymbol{x}-f(\boldsymbol{x})$ is monotone, but differ from our approach as they do not guarantee that the Jacobian of $\boldsymbol{x}-f(\boldsymbol{x})$ is symmetric.
@@ 5.2 Gradients of Sums of (Convex) Ridge Functions
In each of the $2^{n}$ subintervals, $\widetilde{g}_{n,t}$ smoothly interpolates $f$. For a specific $k$ on the RHS of (15), triangle inequality implies
@@
Now, by the Heine-Cantor Theorem [50], $f$ is uniformly continuous on $[0,1]$. Hence let $n\geq n_{0}\in\mathbb{N}$ such that $\forall k,~{}\Delta_{{n_{0}},k}=f\left({k}/{2^{n_{0}}}\right)-f\left({(k-1)}/{2^{n_{0}}}\right)<\epsilon/3$. For each $k$, the second term on the RHS of (17) can be bounded by separately considering the absolute difference between $\tilde{g}(x)$ and $g(x)$ in three different intervals: $i<k$, $i=k$, and $i>k$.
@@ 5.3 Gradients of (convex monotone) transformations of sums of (convex) ridges
Let $\epsilon>0$, and $f,\widehat{f}:[0,1]^{d}\to\mathbb{R}$ be differentiable. If $\sup_{\boldsymbol{x}\in[0,1]^{d}}\|\nabla\widehat{f}(\boldsymbol{x})-\nabla f(\boldsymbol{x})\|\leq\epsilon$ and there exists $\boldsymbol{x}_{0}\in[0,1]^{d}$ with known $f(\boldsymbol{x}_{0})$, then $\widehat{f}$ can recover $f$: $\sup_{\boldsymbol{x}\in[0,1]^{d}}|\widehat{f}(\boldsymbol{x})-f(\boldsymbol{x})|\leq\epsilon\sqrt{d}$
@@
We note that if $\nabla\widehat{f}_{n}$ is a sequence of functions that converges to $\nabla f$, with a corresponding sequence of $\epsilon_{n}>0$ converging to 00, then the recovered sequence of antiderivatives $\widehat{f}_{n}$ converge to $f$ regardless of the dimension $d$. ∎
@@
Let $\mathcal{F}$ be functions in $C^{1}([0,1]^{d})$ expressible as transformations of finite sums of ridge functions. Let $G\in C^{1}([0,1]^{d})$, $g=\nabla_{\boldsymbol{x}}G$, and $\gamma\in C^{0}(\mathbb{R})$. Let $h$ be of the form
@@
where $\beta\in\mathbb{R}$ is a learnable bias. If $\gamma$ universally approximates continuous functions on any compact subset of $\mathbb{R}$ and $g$ universally approximates gradients of sums of ridge functions on $[0,1]^{d}$, then functions of the form $h$ are GradNets and they universally approximate $\nabla\mathcal{F}$.
@@
Let $\mathcal{F}$ be the set of nondecreasing functions in $C^{0}([0,1])$ functions. Let $\tau$ universally approximate nondecreasing functions in $C^{0}([0,1])$ and $\rho$ be uniformly continuous and strictly increasing. Then functions of the form $\rho(\tau(x))$ universally approximate $\mathcal{F}$.
@@
Let $\mathcal{F}$ be functions in $C_{\times}^{1}([0,1]^{d})$ expressible as convex monotone transformations of finite sums of convex ridge functions. Let $h$ be as in (27), where $G\in C^{1}_{\times}([0,1]^{d})$, $\beta\in\mathbb{R}$ is a learnable bias, and $\gamma\in C^{0}(\mathbb{R})$ is nondecreasing and nonnegative. Let $\gamma$ universally approximate nondecreasing and nonnegative functions in $C^{0}(\mathbb{R})$ on compact subset of $\mathbb{R}$. Let $g=\nabla_{\boldsymbol{x}}G$ universally approximate gradients of sums of convex ridge functions on $[0,1]^{d}$. Functions of the form $h$ are mGradNets that universally approximate $\nabla\mathcal{F}$.
@@ 6.1 Modular (Monotone) Gradient Networks (GradNet-M, mGradNet-M)
The design of GradNet-Ms leverages the facts that a linear combinations of GradNets and a composition of a GradNet with a scalar-valued, differentiable function both yield GradNets, as described respectively in Rem. 2 and Thm. 5. The following equations define the GradNet-M and generalize the architecture we previously proposed in [14]:
@@
There are several suitable choices for $\rho_{m},\phi_{m},$ and $\sigma_{m}$ in Cor. 7.1. For example, $\phi_{m}(\boldsymbol{x})=\mathrm{LSE}(\boldsymbol{x})$, $\sigma_{m}(\boldsymbol{x})=\mathrm{softmax}(\boldsymbol{x})$ and $\rho_{m}$ can be any differentiable, monotone nonnegative function on $\mathbb{R}$ such as softplus. As another example, one can take $\sigma_{m}$ to be the elementwise $\mathrm{sigmoid}$ activation function with $\phi_{m}(\boldsymbol{x})=\sum_{i}\mathrm{softplus}(x_{i})$ and $\rho(x)=\alpha x+\beta,\text{ where }\alpha>0$.

### v3

@@ 1 Introduction
Deep neural networks are prized for their ability to parameterize and easily learn complicated, high-dimensional functions. Researchers have devoted substantial effort to developing deep neural networks that achieve state-of-the-art performance on numerous tasks spanning computer vision [1], natural language processing [2], and reinforcement learning [3]. These neural networks are commonly unconstrained and effectively parameterize the space of all functions. Many applications instead require learned functions that exhibit specific properties, which necessitates the design of neural networks corresponding to specific function classes – a problem that has seldom been studied. Constraining neural networks to belong to a particular function class not only enhances interpretability, but also leads to theoretical performance guarantees essential for deploying trained models in safety-critical applications.
In particular, neural networks corresponding to gradients of functions hold significant importance across several science and engineering disciplines. For example, physicists often wish to use a finite set of measurements to characterize the gradient field of a potential function, such as the temperature gradient over a surface. Score-based generative models are another application of learning gradient functions, where neural networks are trained to learn $\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x})$, the score function of an unknown probability distribution [4, 5]. These methods have been applied to image [6, 7] and 3D shape [8] generation.
@@
Contributions: We propose gradient networks (GradNets), neural networks for directly parameterizing gradients of arbitrary functions. We adapt GradNets to monotone gradient networks (mGradNets) that correspond to gradients of convex functions. Our proposed neural network architectures directly model the gradient function $\nabla F$ without first learning the underlying potential function $F$. We also analyze the expressivity of our proposed GradNet and mGradNet networks and prove that they are universal approximators of gradients of general and convex functions, respectively. We further describe methods for modifying these networks to represent popular subsets of these function classes, including gradients of sums of (convex) ridge functions and their (convexity-preserving) transformations. Our theoretical results translate into improved experimental results over existing works. In our gradient field approximation experiments, we find that our architectures achieve up to 15 dB lower mean squared error (MSE) than existing methods. Additionally, we predict the dynamics of the two-body problem and demonstrate an 11 dB improvement over existing methods.
Paper Organization: We first review existing methods for parameterizing and learning gradients of functions in Sec. 2. We then present gradient networks (GradNets) for learning gradients of arbitrary functions in Sec. 3. In Sec. 4, we describe monotone gradient networks (mGradNets): modified GradNets for representing gradients of convex functions. We analyze the expressivity of our proposed GradNets and mGradNets in Sec. 5 and demonstrate methods for composing GradNets and mGradNets to learn (sub)gradients of of the function classes described in Fig. 1. We present architectural modifications in Sec. 6 that empirically yield efficient parameterizations. Finally, in Sec. 7, we evaluate our proposed models on gradient field and Hamiltonian dynamics learning tasks.
@@ 2 Related Work
Several existing works use standard neural networks to directly model the gradient of a potential function [23, 7, 24, 17]. While these methods demonstrate satisfactory empirical performance, they lack theoretical justification and are often considered heuristic approaches. It is not guaranteed, and in fact highly unlikely, that an arbitrary neural network with matching input and output dimension corresponds to the gradient of a scalar-valued function. More formally, let $\mathcal{F}$ be a set such that each function $F\in\mathcal{F}$ is the gradient of a scalar-valued function. Given an $F\in\mathcal{F}$, standard neural network architectures, e.g., multilayer perceptrons (MLPs) and convolutional neural networks (CNNs), with nonpolynomial activations, can be used to approximate $F$. However, these architectures offer no guarantee that the learned function $\widehat{F}$ is itself in $\mathcal{F}$. In contrast, the methods and network architectures discussed in this paper ensure that the approximator $\widehat{F}$ is always a member of $\mathcal{F}$, thereby enhancing interpretability and enabling robust theoretical guarantees in practice.
@@
Rather than directly approximating the gradient of a function using a standard neural network, an alternative approach may be to first parameterize and learn the underlying scalar potential function using a standard neural network. Subsequently differentiating the network with respect to its input yields the desired gradient. However, closely approximating a target function $F$ with an arbitrary approximator $G$ does not guarantee that $\nabla G$ closely approximates $\nabla F$ [25]. Empirical evidence in [26] confirms this fact when the approximating functions are neural networks. Other works directly train the gradient map of a neural network that parameterizes a potential function. For instance, [21] considers the potential $F(\boldsymbol{x})=\frac{1}{2}\|\boldsymbol{x}-f_{\boldsymbol{\theta}}(\boldsymbol{x})\|_{2}^{2}$ with neural network $f_{\boldsymbol{\theta}}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and uses automatic differentiation to obtain a gradient network architecture $\nabla F(\boldsymbol{x})$. Nevertheless, these approaches often exhibit ill-behaved product structures in practice [27] and, due to the Runge phenomena [28], optimizing their parameters becomes cumbersome as the input dimension grows [29].
@@
Feedforward neural networks have also been proposed to directly parameterize gradients of specific classes of potential functions, namely potentials $F(\boldsymbol{x})$ expressible as the sum of ridge functions: $F(\boldsymbol{x})=\sum_{i=1}^{n}\psi_{i}(\boldsymbol{w}_{i}^{\top}\boldsymbol{x})$. Such potentials were considered in the fields of experts framework proposed in [30]. The fields of experts framework has inspired several methods for learning image priors that effectively generalize the “transform domain thresholding” approach discussed in [17, 31]. These include works like [32, 33] that correspond to nonconvex potentials and respectively achieve commendable performance for image restoration and magnetic resonance imaging reconstruction. More recently, [34, 35] propose neural networks with elementwise, learnable spline activations to learn gradients of potentials expressible as sums of ridge functions (class G in Fig. 1). The networks can be modified to correspond to gradients of sums of convex ridge functions (class H in Fig. 1) by restricting the splines to be nondecreasing. In this paper, we consider parameterizing the activations as learnable neural networks and discuss how these networks can be adapted to represent (sub)gradients of convex potentials. Furthermore, the class of functions expressible as a sum of convex ridges, as in [34, 35], does not include several common convex functions including
@@ 3 Gradient Networks (GradNet)
Prop. 0.1 states that (3) is a GradNet if the activation function $\sigma$ has an antiderivative. It hence permits the use of group activations such as softmax. Furthermore, to guarantee existence of an antiderivative, it is sufficient, but not necessary, that $\sigma$ is differentiable and its Jacobian is symmetric everywhere. For example, the elementwise activation $\mathrm{ReLU}(\boldsymbol{x})=\max(0,\boldsymbol{x})$ is continuous and nondifferentiable, but it is the gradient of $\sum_{i}\max(0,\frac{1}{2}x_{i}^{2})$. Architectures of the form in (3), with a continuous elementwise activation function $\sigma(\boldsymbol{x})=\begin{bmatrix}\sigma_{1}(x_{1})\ \dots\ \sigma_{m}(x_{m})\end{bmatrix}^{\top}$, include “transform domain thresholding” methods [17, 31] where the activation is a continuous thresholding operator. Similar architectures appear in [35], which uses linear spline activations, and in [33], which parameterizes each $\sigma_{i}$ as a linear combination of Gaussian radial basis functions. An alternative approach is to use neural networks to parameterize the elementwise activation functions.
@@ 4 Monotone Gradient Networks (mGradNet)
Rem. 3 follows from Def. 1 and Remark 2.2 in [35], which implies monotonicity of $g$. It demonstrates how to construct an mGradNet using a GradNet with known Lipschitz constant. Rem. 3 relates to several Regularization by Denoising (RED) works that design learnable functions of the form $\boldsymbol{x}-f(\boldsymbol{x})$ where $f(\boldsymbol{x})$ is constrained to be nonexpansive [45, 46]. These methods differ from our approach as they do not guarantee that the Jacobian of $\boldsymbol{x}-f(\boldsymbol{x})$ is symmetric.
@@ 5.2 Gradients of Sums of (Convex) Ridge Functions
In each of the $2^{n}$ subintervals, $\widetilde{g}_{n,t}$ interpolates $f$. For a specific $k$ on the RHS of (15), triangle inequality implies
@@
Now, by the Heine-Cantor Theorem [50], $f$ is uniformly continuous on $[0,1]$. Hence, let $n\geq n_{0}\in\mathbb{N}$ such that $\forall k,~{}\Delta_{{n_{0}},k}=f\left({k}/{2^{n_{0}}}\right)-f\left({(k-1)}/{2^{n_{0}}}\right)<\epsilon/3$. For each $k$, the second term on the RHS of (17) can be bounded by separately considering the absolute difference between $\tilde{g}(x)$ and $g(x)$ in three different intervals: $i<k$, $i=k$, and $i>k$.
@@ 5.3 Gradients of (convex monotone) transformations of sums of (convex) ridges
Let $\epsilon>0$, and $f,\widehat{f}:[0,1]^{d}\to\mathbb{R}$ be differentiable. If $\sup_{\boldsymbol{x}\in[0,1]^{d}}\|\nabla\widehat{f}(\boldsymbol{x})-\nabla f(\boldsymbol{x})\|\leq\epsilon$ and there exists $\boldsymbol{x}_{0}\in[0,1]^{d}$ with known $f(\boldsymbol{x}_{0})$, then $\widehat{f}$ can approximate $f$: $\sup_{\boldsymbol{x}\in[0,1]^{d}}|\widehat{f}(\boldsymbol{x})-f(\boldsymbol{x})|\leq\epsilon\sqrt{d}$
@@
We note that if $\nabla\widehat{f}_{n}$ is a sequence of functions that converges to $\nabla f$, with a corresponding sequence of $\epsilon_{n}>0$ converging to 00, then the sequence of antiderivatives $\widehat{f}_{n}$ converges to $f$ regardless of the dimension $d$. ∎
@@
Let $\mathcal{F}$ be functions in $C^{1}([0,1]^{d})$ expressible as transformations of finite sums of ridge functions. Let $G\in C^{1}([0,1]^{d})$, $g(\boldsymbol{x})=\nabla_{\boldsymbol{x}}G(\boldsymbol{x})$, and $\gamma\in C^{0}(\mathbb{R})$. Let $h$ be of the form
@@
where $\beta\in\mathbb{R}$ is a learnable bias. If $\gamma$ universally approximates continuous functions on any compact subset of $\mathbb{R}$ and $g$ universally approximates gradients of sums of ridge functions on $[0,1]^{d}$, then functions of the form $h$ are GradNets that universally approximate $\nabla\mathcal{F}$.
@@
Let $\mathcal{F}$ be the set of nondecreasing functions in $C^{0}([0,1])$. Let $\tau$ universally approximate nondecreasing functions in $C^{0}([0,1])$ and $\rho$ be uniformly continuous and strictly increasing. Then functions of the form $\rho(\tau(x))$ universally approximate $\mathcal{F}$.
@@
Let $\mathcal{F}$ be functions in $C_{\times}^{1}([0,1]^{d})$ expressible as convex monotone transformations of finite sums of convex ridge functions. Let $h$ be as in (27), where $G\in C^{1}_{\times}([0,1]^{d})$, $\beta\in\mathbb{R}$ is a learnable bias, and $\gamma\in C^{0}(\mathbb{R})$ is nondecreasing and nonnegative. Let $\gamma$ universally approximate nondecreasing and nonnegative functions in $C^{0}(\mathbb{R})$ on compact subset of $\mathbb{R}$. Let $g(\boldsymbol{x})=\nabla_{\boldsymbol{x}}G(\boldsymbol{x})$ universally approximate gradients of sums of convex ridge functions on $[0,1]^{d}$. Functions of the form $h$ are mGradNets that universally approximate $\nabla\mathcal{F}$.
@@ 6.1 Modular (Monotone) Gradient Networks (GradNet-M, mGradNet-M)
The design of GradNet-Ms leverages the facts that a linear combination of GradNets and a composition of a GradNet with a scalar-valued, differentiable function both yield GradNets, as described respectively in Rem. 2 and Thm. 5. The following equations define the GradNet-M and generalize the architecture we previously proposed in [14]:
@@
There are several suitable choices for $\rho_{m},\phi_{m},$ and $\sigma_{m}$ in Cor. 7.1. For example: $\phi_{m}(\boldsymbol{x})=\mathrm{LSE}(\boldsymbol{x})$, $\sigma_{m}(\boldsymbol{x})=\mathrm{softmax}(\boldsymbol{x})$ and $\rho_{m}$ any differentiable, monotone nonnegative function on $\mathbb{R}$ such as softplus. As another example, one can take $\sigma_{m}$ to be the elementwise $\mathrm{sigmoid}$ activation function with $\phi_{m}(\boldsymbol{x})=\sum_{i}\mathrm{softplus}(x_{i})$ and $\rho(x)=\alpha x+\beta,\text{ where }\alpha>0$.